Determine the force between two positively charged particles that are separated by a distance of 3.71cm and that have charges of 1.4 x 10-5C and 3.2 x 10-5C.



1
Expert's answer
2022-06-14T08:41:14-0400
"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{1.4*10^{-5}*3.2*10^{-5}}{0.0371^2}=2900\:\rm N"